OpenStudy (anonymous):

I will medal and fan, Eva makes a few dozen cupcakes and sells them for $8 a dozen. Which function represents the relationship between the money Eva earns, m, and the number of cupcakes she sells, c?

OpenStudy (anonymous):

@Pawanyadav

OpenStudy (anonymous):

whats up

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

To convert a total number of cupckaes to 'number of dozens' we first dvide by 12 i.e. \[D= \frac{c}{12}\] ex: \[\frac{24 cupcakes}{12} = 2 dozen\]we can then multiply this by the price per dozen \[m=8D\] and then substitute \[m=8 \times \frac{c}{12}\] or \[m=\frac{8c}{12}=\frac{2c}{3}\]

OpenStudy (anonymous):

f(c) = c × 8 f(m) = c × 8 f(c) = m × 8 f(m) = m × 8 f(m) = c × m which on is it @twvogels